OpenStudy (anonymous):

A complex number Z satisfies |Z-(-2+2sqrt(3)i)|<=2

OpenStudy (anonymous):

Find the least psosible values of |Z| so Z is going to be a circle (x+2)^2+(y-2sqrt(3))^2<=2

OpenStudy (anonymous):

(x+2)^2+(y-2sqrt(3))^2<=2^2 * how do i find the point clsoest to origin?

OpenStudy (sirm3d):

algebra-wise, this point (x,y) lies in the intersection of the line through the origin and \[\left( -2,2\sqrt{3} \right)\] and the circle of radius and and center at \[\left( -2,2\sqrt{3} \right)\]

OpenStudy (sirm3d):

the radius of the circle is two

OpenStudy (anonymous):

hm why is the shortest from the origin to centre of circle?

OpenStudy (anonymous):

hi?

OpenStudy (sirm3d):

i didn't get the "why" question

OpenStudy (anonymous):

what you saying is that the point Z closest to the Origin is the point of intersection of the line from origin to centre of circle right

OpenStudy (sirm3d):

right

OpenStudy (sirm3d):

one point of intersection is nearest the origin, the other is farthest

OpenStudy (anonymous):

why what is your reasoning for that?

OpenStudy (anonymous):

HII REPLY

OpenStudy (sirm3d):

that's a fact from analytic geometry, unless you haven't taken up that course

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

im in high school!

OpenStudy (sirm3d):

how about geometry? not yet?

OpenStudy (anonymous):

im not sure.. can you do a quick show?

OpenStudy (sirm3d):

 will this suffice?

OpenStudy (anonymous):

hmm yeah

OpenStudy (anonymous):

oh so.. shortest dist between a point and a circle is always going to be line perpendicular to tangent at any point

OpenStudy (anonymous):

or something like that..

OpenStudy (sirm3d):

yes

OpenStudy (anonymous):

what about find the greatest posible value of Arg(Z)